A researcher uses two samples, each with n = 15 participants, to evaluate the mean difference in performance scores between 8 please explain how to do this because i dont get it thank you
Answer #1

Ans:

Right tailed test.

(as claim is that older children will have higher scores)

pooled variance=30

standard error for difference=sqrt(30*((1/15)+(1/15)))=2

test statistic:

t=5/2

t=2.5

df=15+15-2=28

p-value=tdist(2.5,28,1)=0.0093

As,p-value is less than 0.05 and 0.01,we can reject the null hypothesis at both alpha=0.05 or 0.01

Option b is correct.

Reject the null hypothesis with either alpha=0.05 or alpha=0.01